### v1

# On Biologically Plausible Learning in Continuous Time

Abstract
Biological learning unfolds continuously in time, yet most algorithmic models rely on discrete updates and separate inference and learning phases. We study a continuous-time neural model that unifies several biologically plausible learning algorithms and removes the need for phase separation. Rules including stochastic gradient descent (SGD), feedback alignment (FA), direct feedback alignment (DFA), and Kolen–Pollack (KP) emerge naturally as limiting cases of the dynamics. Simulations show that these continuous-time networks stably learn at biological timescales, even under temporal mismatches and integration noise. Through analysis and simulation, we show that learning depends on temporal overlap: a synapse updates correctly only when its input and the corresponding error signal coincide in time. When inputs are held constant, learning strength declines linearly as the delay between input and error approaches the stimulus duration, explaining observed robustness and failure across network depths. Critically, robust learning requires the synaptic plasticity timescale to exceed the stimulus duration by one to two orders of magnitude. For typical cortical stimuli (tens of milliseconds), this places the functional plasticity window in the few-second range, a testable prediction that identifies seconds-scale eligibility traces as necessary for error-driven learning in biological circuits.
1 Introduction
Understanding how biological circuits learn has long been a central challenge at the interface of neuroscience and machine learning. Many algorithmic proposals for biologically plausible learning have been put forward, including feedback alignment (FA; Lillicrap et al. (2016)), direct feedback alignment (DFA; Nøkland (2016)), and Kolen–Pollack / weight mirrors (KP; Kolen and Pollack (1994); Akrout et al. (2019)). These methods attempt to relax the strong requirements of backpropagation, such as exact weight transport, while preserving the ability to optimize deep networks. Such rules are attractive both as potential models of cortical learning and as candidates for training in neuromorphic or analog hardware.
Two key aspects of real biological systems, however, are conspicuously absent from most of these formulations: learning in biology unfolds continuously in time, and does not rely on separate learning and inference phases. Realistic neurons operate with finite conduction and integration times, so neither inference nor plasticity can be assumed to occur instantaneously or in synchronized steps (Kandel et al., 2000). By contrast, existing algorithmic models are cast in discrete steps that alternate between inference and learning phases, effectively assuming that forward and backward signals are globally and instantaneously synchronized (Whittington and Bogacz, 2019). Therefore, the existing biologically plausible algorithms fail to answer a basic question: can learning be achieved in continuous time, without artificial phase separation, and at biologically realistic timescales?
In this work, we address this gap by formulating a continuous-time neural model in which both neural states and synaptic weights evolve according to coupled first-order differential equations. Each neuron receives a feedforward drive and a modulatory error drive, and updates its synapses through a local two-signal rule. Crucially, inference and learning occur simultaneously: there is no global clock and no need for phase separation as all state variables evolve under the same ODE. The model is governed by distinct time constants for fast signal propagation, intermediate plasticity, and slow synaptic decay, an ordering that mirrors the hierarchy of timescales in real neural tissue.
Our contributions are threefold:
  1. 1.
    Continuous-time formulation. We introduce a dynamical-systems model where inference and learning unfold simultaneously, without phase separation. Classical rules and biologically plausible algorithms such as SGD, FA, DFA, and KP emerge as limiting cases under timescale separation (Section 3).
  2. 2.
    Temporal learning analysis. We analyze the effect of timing mismatches and show that learning is determined by the temporal overlap between input and error. This predicts failure when delay approaches the sample duration, depth fragility from accumulated lags, and robustness patterns observed in simulation (Sections 45).
  3. 3.
    Biological timescales. We demonstrate effective learning at biologically realistic timescales and derive a quantitative constraint: plasticity windows must outlast the stimulus by an order of magnitude, placing the functional eligibility regime in the few-second range (Section 5.3).
Together, these results support the biological feasibility of rules like KP, FA, and DFA when cast in continuous time. They show that plasticity arises from temporal overlap between input and error, explaining the robustness of these algorithms and providing a physically grounded path toward continuous-time learning in both biological and artificial systems.
2 Related Work
Related families of algorithms, such as contrastive Hebbian learning (Xie and Seung, 2003), equilibrium propagation (Scellier and Bengio, 2017), and other energy-based formulations (Hopfield, 1982; Bengio and Fischer, 2015), also instantiate continuous-time learning rules. These approaches rely on symmetric connectivity or two-phase equilibrium dynamics, which differ from the heterosynaptic two-signal rules that are our focus here. While our model shares the spirit of framing learning as a dynamical process, we restrict our focus in this work to error propagation style rules (FA, DFA, KP) and their continuous-time realizations.
Both neural ODEs (Chen et al., 2018) and our neural differential-equation model cast network computation as a continuous-time dynamical system, replacing discrete layers/updates with ODE flows over time. In standard neural ODEs, parameters are fixed during the forward solve, and gradients are typically recovered by integrating an adjoint ODE backward in time (Chen et al., 2018) or by differentiating through the solver (Baydin et al., 2018). By contrast, our model couples learning and inference in one forward-in-time system: both neural states and parameters evolve by ODEs, with weights updated online via locally computed and propagated error terms.
Concurrently, “self-assembling” heterosynaptic circuits specifying four plastic synapses linking forward and feedback streams have been shown to match backpropagation on standard benchmarks (Liao et al., 2024). A key insight from this line of work is that error-propagation rules such as feedback-style algorithms (FA, DFA, KP) can emerge without requiring explicit weight symmetry, and only using local rules. This further motivates our focus on these algorithms as being attractive candidates for biologically plausible and hardware-realizable learning.
Biological relevance of feedback-type algorithms has been argued in various works. For example, Lillicrap et al. (2020) argued how FA may be realizable in the cerebellum, where the Purkinje cells are the trainable neurons, and the one-to-one climbing fibres carry in the feedback error signals. Koplow et al. (2025) computationally showed how these algorithms are consistent with the widely observed Hebbian and anti-Hebbian plasiticities in the brain. However, it is not yet clear whether these algorithm can work in the continuous-time, and if so, whether they work at biologically reasonable timescales. This is a key question we address.
3 A Continuous-Time Neural Network Model
[…]
Figure 1: Neuron design with two archetypal topologies. At left (zoom): continuous-time heterosynaptic neuron model whose weights evolve according to plasticity rules. Neurons receive forward input x and produce activated output $z$ with weights w. Error signals e enter through modulatory weights v and drive plasticity of w. In center: layerwise error propagation topology, where errors are propagated to previous layers via backward weights. At right: direct error propagation topology, where error signals are broadcast to previous layers directly.
Figure 1 illustrates the continuous-time heterosynaptic neuronal systems we study. The forward pathway, parameterized by $\mathbf{w}$ and with activation $\sigma$, maps inputs $\mathbf{x}$ to neuronal output $z$. The error pathway, parameterized by $\mathbf{v}$, receives an error signal $\mathbf{e}$ and provides a modulatory influence on learning. Learning is heterosynaptic: the forward weights $\mathbf{w}$ update in proportion to the modulatory drive $(\mathbf{v}^{\top}\mathbf{e})$, while the error weights $\mathbf{v}$ update in proportion to the forward drive $(\mathbf{w}^{\top}\mathbf{x})$. Let layer $l$ have width $d_{l}$. Stacking the per-neuron updates, we can use the following notation: $\mathbf{z}_{l-1}(t)\in\mathbb{R}^{d_{l-1}}$, $\mathbf{z}_{l}(t)\in\mathbb{R}^{d_{l}}$, $W_{l}\in\mathbb{R}^{d_{l}\times d_{l-1}}$, and error source $e_{l}^{\text{src}}(t)$. Define
$\dot{\mathbf{z}}_{l}=\frac{-\mathbf{z}_{l}+\sigma_{l}(W_{l}^{\top}\mathbf{z}_{l-1})}{\tau_{\mathrm{prop}}},\quad\dot{W}_{l}=-\frac{W_{l}}{\tau_{\mathrm{dec}}^{W}}+\frac{\mathbf{z}_{l-1}\,(V_{l}^{\top}\bm{\epsilon}_{l})^{\top}}{\tau_{\mathrm{plas}}^{W}},\quad\dot{V}_{l}=-\frac{V_{l}}{\tau_{\mathrm{dec}}^{V}}+\frac{(W_{l}^{\top}\mathbf{z}_{l-1})\,\bm{\epsilon}_{l}^{\top}}{\tau_{\mathrm{plas}}^{V}}.$ (1)
Here $\tau_{\mathrm{prop}}$ is the neuronal propagation time constant, setting how quickly $\mathbf{z}_{l}$ relaxes to its driven input. The constants $\tau_{\mathrm{plas}}^{W},\tau_{\mathrm{plas}}^{V}$ control how rapidly synaptic plasticity occurs when presynaptic and modulatory drives coincide. The constants $\tau_{\mathrm{dec}}^{W},\tau_{\mathrm{dec}}^{V}$ govern passive weight decay, setting the forgetting timescale. For the biological interpretation of these timescales, see Section 5.3. These equations arise by stacking per-neuron heterosynaptic updates; see Appendix A for the per-neuron update rules.
We emphasize that these equations describe rate-based dynamics, not spiking dynamics: the neuronal state $z(t)$ evolves continuously without threshold-triggered resets or discrete spikes (Hopfield, 1984; Wilson and Cowan, 1972). This abstraction allows us to focus on continuous-time error-propagation rules and makes the analysis tractable. While the lack of spikes is a limitation that excludes important timing effects and spike-based coding strategies, it can still serve as a reasonable model for certain classes of cortical neurons whose average firing rates are well-approximated by continuous variables (Wilson and Cowan, 1972). Moreover, rate-based neurons align naturally with the error-propagation algorithms we study (FA, DFA, KP), which are themselves formulated in terms of continuous activations rather than discrete events.
These dynamics can be understood as a continuous-time realization of the heterosynaptic two-signal principle. Under separation of timescales, we may treat neuronal activity as equilibrated relative to the slower synaptic updates. In this quasi-stationary regime, the fast variables satisfy $\mathbf{z}_{l}=\sigma(W_{l}^{\top}\mathbf{z}_{l-1})$, and the weight dynamics reduce to the heterosynaptic updates analyzed in Ziyin et al. (2025), where local stability guarantees each neuron implements a gradient step and global consistency ensures alignment across layers. At full stationarity ($\dot{\mathbf{z}}=\dot{W}=\dot{V}=0$) and with equal size, the fixed-point equations imply $W_{l}\propto\mathbf{z}_{l-1}(V_{l}^{\top}\bm{\epsilon}_{l})^{\top}$ and $V_{l}\propto(W_{l}^{\top}\mathbf{z}_{l-1})\bm{\epsilon}_{l}^{\top}$, so that $W_{l}$ converges to a configuration consistent with $V_{l}^{\top}$. In this limit, the effective update to $W_{l}$ is the outer product of the input with the backpropagated error at that layer—recovering the canonical backpropagation rule as the stationary solution of the dynamics.
The center and right panels of Figure 1 show how these neurons connect into multilayer networks. Under timescale separation $\tau_{\mathrm{prop}}{}\ll\tau_{\mathrm{plas}}\ll\tau_{\mathrm{dec}}$, the activation states equilibrate quickly, and weights update quasi-statically. Outputs track their steady state $\mathbf{z}_{l}\approx\sigma_{l}(W_{l}^{\top}\mathbf{z}_{l-1})$ within a sample window. Averaging the weight ODE over one input data presentation gives a weight update $\Delta W_{l}\propto\mathbb{E}_{\text{window}}\left[\mathbf{z}_{l-1}(V_{l}^{\top}\bm{\epsilon}_{l})^{\top}\right]$. The effective weight update rule corresponding to different learning algorithms depends only on the definitions of the feedback weights $V_{l}$ and error drive $\bm{\epsilon}_{l}$. Given the error $\mathbf{e}_{L}$ as the gradient of the loss function with respect to the output layer:
SGD. Under the constraint $V_{l}=W_{l}^{\top}$, the error signals are equal to the true backpropagated error, $\bm{\epsilon}_{l}=\mathbf{e}_{l}$. Then, $\Delta W_{l}\propto\mathbb{E}[\mathbf{z}_{l-1}\mathbf{e}_{l}^{\top}]$, which is exactly the gradient of the loss.
Feedback Alignment (FA). Set $V_{l}$ to fixed random matrices that project the error at each layer backwards to the previous layer. Then, $\Delta W_{l}\propto\mathbb{E}[\mathbf{z}_{l-1}(V_{l}^{\top}\bm{\epsilon}_{l})^{\top}]$, which is the standard FA update. Gradient alignment arises empirically (Lillicrap et al., 2016).
Direct Feedback Alignment (DFA). Set $V_{l}$ to fixed random matrices. Broadcast the output-layer error $\mathbf{e}_{L}$ to all hidden layers, such that $\bm{\epsilon}_{l}=\mathbf{e}_{L}$. Then, $\Delta W_{l}\propto\mathbb{E}[\mathbf{z}_{l-1}(V_{l}^{\top}\mathbf{e}_{L})^{\top}]$, which is the DFA update. Each layer learns from a direct projection of the global error (Nøkland, 2016).
Kolen-Pollack (KP) / Weight-Mirror Methods. Allow $V_{l}$ to evolve under the full plasticity rules in equation 1. $V_{l}$ tracks the correlation between the forward drive $W_{l}^{\top}\mathbf{z}_{l-1}$ and the error $\bm{\epsilon}_{l}$. In expectation, this update combined with weight decay pushes $V_{l}$ towards $W_{l}^{\top}$. Thus KP can be viewed as a dynamical mechanism that drives $V_{l}\to W_{l}^{\top}$, in contrast to SGD which assumes this equality is enforced from the start (Akrout et al., 2019; Kolen and Pollack, 1994).
We simulate training and inference of these models using ODE solvers (Kidger, 2021). Inputs to the network are driven by the dataset’s input signals, and the error at the output neurons in layer $L$ are $\mathbf{e}=\partial\mathcal{L}/\partial\mathbf{z}_{L}$. Figure 3 shows an example of the output-layer dynamics of early training. Each data point is presented to the network’s input neurons for a fixed duration of time (sample time), with interpolation between the images done smoothly over a constant, much shorter duration (buffer time). Rather than impose the corresponding constraints and clamps on weights as traditional SGD/KP/FA/DFA prescribe, our experiments leave both $W$ and $V$ free to learn. We find that this makes our network architecture more general without sacrificing task performance.
Figure 2: Zoomed-in view of output neuron dynamics during early training. The network input and corresponding error term changes during the buffer time periods. Over a single sample, the neuron outputs move only minutely.
Figure 3: Sample train/test accuracy during training. Training set accuracy is measured as the moving average of instantaneous correct/incorrect predictions on the network input signals. Test set accuracy is measured by evaluating the network with frozen weight dynamics.
During evaluation, parameters are held fixed and plasticity and decay dynamics are frozen: $\dot{W}=\dot{V}=0$ and $\bm{\epsilon}\equiv 0$. Only inference dynamics $\dot{\mathbf{z}}$ are active. The model’s prediction is read out at the end of the input presentation window, immediately before the next sample is shown.
4 Timing Robustness
A central question is: under what conditions does a synapse receive a correct update? Two key considerations are (i) the temporal mismatch between input and error signals, and (ii) the rate at which the input changes. Figure 4 illustrates the dynamics of a final-layer weight when the error signal arrives earlier or later than the corresponding input. During the mismatch period, the instantaneous weight update is incorrect, leading to a cumulative update that deviates from the non-delayed case.
Figure 4: Single-neuron dynamics with different relative timings of the error and input signals. Left: error signal arrives early. Right: error signal is delayed. The bottom panel of both plots depicts the cumulative weight change and shows that in the presence of delay, the weight accumulates a biased gradient update compared to the case where there was no error delay.
To analyze robustness, we isolate only the part of the weight update that is informative: that which is proportional to the correlation between presynaptic activity and the matching error drive. Vectorized, the informative weight change accumulated over a presentation window of length $T$ is
$\Delta W_{l}\;\propto\;\int_{0}^{T}\mathbf{z}_{l-1}(t)\,\big(V_{l}^{\top}\bm{\epsilon}_{l}(t)\big)^{\top}\,k_{\tau_{\mathrm{plas}}}(t)\;dt,\qquad k_{\tau_{\mathrm{plas}}}(t)=\exp\!\big(-(T-t)/\tau_{\mathrm{plas}}\big),$ (2)
where the causal exponential kernel arises from the low-pass plasticity dynamics and weights more recent coincidence more strongly. For a single synapse $(i\to j)$, this reduces to the scalar form
$\Delta(W_{l})_{ij}\;\propto\;\int_{0}^{T}(\mathbf{z}_{l-1}(t))_{i}\;\big((V_{l})_{:,j}^{\top}\bm{\epsilon}_{l}(t)\big)\;k_{\tau_{\mathrm{plas}}}(t)\;dt.$ (3)
This makes clear that learning depends on the temporal cross-correlation between the presynaptic drive $(\mathbf{z}_{l-1})_{i}$ and the local modulatory/error drive at neuron $j$.
Piecewise-constant inputs with delay. Assume $(\mathbf{z}_{l-1})_{i}(t)$ is active on $[0,T]$ and the error drive is active on $\Delta,\Delta+T]$ (same duration, delayed by $\Delta$). In the fast-propagation limit $\tau_{\mathrm{prop}}{}\ll\tau_{\mathrm{plas}}$, the expected update becomes
$\mathbb{E}[\Delta(W_{l})_{ij}]$ $\propto\int_{t_{0}}^{t_{1}}\exp\!\big((t-T)/\tau_{\mathrm{plas}}\big)\,dt$
$=\tau_{\mathrm{plas}}\!\left(e^{-(T-t_{1})/\tau_{\mathrm{plas}}}-e^{-(T-t_{0})/\tau_{\mathrm{plas}}}\right)=\tau_{\mathrm{plas}}\,e^{-(T-t_{1})/\tau_{\mathrm{plas}}}\!\left(1-e^{-L/\tau_{\mathrm{plas}}}\right),$ (4)
where $t_{0}=\max(0,\Delta)$, $t_{1}=\min(T,\Delta+T)$, and $L=t_{1}-t_{0}=(T-|\Delta|)_{+}$.
Flat-kernel limit and prediction. When $T\ll\tau_{\mathrm{plas}}$ (the regime standard in our experiments), $k_{\tau_{\mathrm{plas}}}$ is approximately constant over $[0,T]$, so equation 4 reduces to the symmetric triangular law
$\mathbb{E}[\Delta(W_{l})_{ij}]\;\propto\;(T-|\Delta|)_{+}.$ (5)
Thus learning succeeds if and only if input and error overlap in time, and it degrades sharply as $|\Delta|\to T$. When $T$ approaches $\tau_{\mathrm{plas}}$, the exact expression equation 4 predicts a mildly skewed triangle that up-weights late-arriving errors (positive $\Delta$) relative to equally early ones; the skew vanishes continuously as $\tau_{\mathrm{plas}}/T\to\infty$.
A fixed overlap budget. For delay $\Delta$, define the correct-overlap set $C(\Delta)=[0,T]\cap[\Delta,\Delta+T]$ of length $L=(T-|\Delta|)_{+}$ and mismatched set $I(\Delta)=[0,T]\setminus C(\Delta)$. With plasticity kernel $k(t)$, let
$K_{C}(\Delta)=\!\!\int_{C(\Delta)}\!k(t)\,dt,\quad K_{I}(\Delta)=\!\!\int_{I(\Delta)}\!k(t)\,dt,$
so that $K_{C}(\Delta)+K_{I}(\Delta)=K_{T}:=\int_{0}^{T}k(t)\,dt$. Thus overlap and mismatch trade off under a fixed budget. In the flat-kernel case $k\equiv 1$, this reduces to L+(TL)=TL+(T-L)=T, yielding $\mathbb{E}[\Delta(W_{l})_{ij}]\propto L=(T-|\Delta|)_{+}$, i.e. accuracy improves with either larger $T$ or smaller $|\Delta|$.
Our analysis predicts and experiments confirm near-symmetry between early and late error in the flat-kernel regime. By contrast, biological plasticity is causally gated: synaptic activity first writes a short-lived eligibility trace, and only subsequently arriving modulatory/error signals consolidate it into weight change (Yagishita et al., 2014). Thus, early error fails to drive learning even though late error can—an asymmetry our simplified model compresses into a symmetric overlap law. Incorporating an explicit eligibility gate would recover this asymmetry without altering our conclusions about temporal overlap and timescale separation; we leave such extensions for future work.
5 Experiments
Running these experiments is computationally demanding. Each training and evaluation run requires explicit forward integration of a high-dimensional, stiff system of coupled ODEs. Neuronal activations evolve on millisecond timescales, while meaningful learning unfolds over hundreds to thousands of seconds. This range of time scales spans six orders of magnitude, forcing solvers to take very small steps across very long horizons.
We implement our simulations in JAX (Bradbury et al., 2018) and use the Diffrax library (Kidger, 2021) for ODE integration. Although we do not rely on JAX’s autodifferentiation, its just-in-time compilation and efficient array operations provide substantial performance gains for these large-scale simulations. For integration we use Tsit5 (Tsitouras, 2011), an explicit 5th-order Runge–Kutta method with an embedded 4th-order error estimate. This combination is accurate and stable without resorting to computationally expensive implicit methods. The embedded estimate enables adaptive step sizing under PID control, which we find essential for efficiently handling stiffness and long simulation horizons.
Some of the parameter sweeps reported here, particularly the heatmaps over error delay and sample duration, require over 500 compute hours to complete. This reflects the intrinsic difficulty of simulating continuous-time learning rules, where propagation, plasticity, and decay processes must be faithfully resolved simultaneously. These results represent a substantial computational effort and demonstrate that continuous-time neural networks can be simulated effectively despite these challenges.
In this section, we evaluate three experimental regimes. Section 5.1 demonstrates that networks can robustly learn so long as error signals overlap with inputs. Section 5.2 demonstrate that deeper networks accumulate propagation lag, making them less tolerant to error delays and requiring longer sample times for stable learning. Finally, Section 5.3 shows that our framework operates effectively under synaptic, plasticity, and decay timescales that align with known cortical physiology.
5.1 Direct Error Routing
The direct error routing topology sets the local error at each layer to be the global error (from the output layer), $\bm{\epsilon}_{l}=\mathbf{e}_{L}$. This corresponds to a Direct Feedback Alignment scheme where the backward weights $V_{l}$ are also learned. Figure 5 examines the accuracy of the direct error routing topology on the $7\times 7$ downsampled MNIST dataset as the error-signal delay and sample duration are varied.
Figure 5: Evaluation of the direct error routing topology on the $7\times 7$ downsampled MNIST dataset. The $x$-axis denotes the temporal delay between input signal and label: negative values indicate the label arrives before the input, while positive values indicate the label arrives after the input. At left is a network with 1 hidden layer of 49 neurons, and at right is a network with 2 hidden layers of 49 and 32 neurons. The learning is fairly robust until the delay exceeds the sample time.
In the direct error routing topology, each layer receives a direct copy of the error signal. As a result, all neurons in the network receive the error signal with a propagation delay of one synaptic length. Learning fails when the error signal delay is approximately equal to or longer than the sample duration, since the error signal has zero overlap with its corresponding input. In these regions (accuracy $\sim 10\%$), weights only change due to spurious correlations between mismatched input-label pairs. Transitional regions ($\sim 70\%$ accuracy) arise when error and input partially overlap: some useful updates occur, but they are mixed with incorrect updates from mismatched periods. These networks have a fixed $\tau_{\mathrm{plas}}=10$s while the sample time $T$ lies in 0.01-0.25s, putting them primarily in the flat-kernel learning limit. As a result, learning responds fairly symmetrically to both early and late error signals.
5.2 Layerwise Error Routing
The layerwise error routing topology propagates error from the output layer step by step to previous layers via $V_{l}^{\top}\bm{\epsilon}_{l}$. This corresponds to a KP or weight mirroring scheme where the backward weights $V_{l}$ are learned and converge to transposes of the forward weights. Figure 6 examines the accuracy of the layerwise error routing topology on a synthetic two-dimensional circles dataset. This task is more nonlinear than the $7\times 7$ downsampled MNIST task, and requires deeper ReLU networks to represent an accurate decision boundary.
Figure 6: Evaluation of the layerwise error routing topology on the circle dataset as error signal delay and sample duration are swept. The network topologies increase in the number of hidden layers, where each added hidden layer has 24 neurons. From left to right, the networks have one hidden layer of 24 ReLU neurons, two hidden layers of 24 ReLU neurons, and three hidden layers of 24 ReLU neurons. Layerwise error routing imposes stricter requirements on delay due to longer error propagation paths.
Unlike the direct error routing topology, the layerwise error routing topology suffers more from finite propagation times: not only does the input signal have to propagate to the output layer, but the error signal must also propagate backward from the output layer back to the input layer. The heatmaps show that these deeper networks require longer sample times and are more sensitive to delay in the error signal. Finite propagation speed means that deeper networks accumulate lag, degrading the learning signal especially in early layers. Longer sample durations mitigate this effect by increasing the window where inputs and their true errors coincide. In the language of Section 4, this increases $T$ relative to a fixed $\Delta$, resulting in a greater correct update $\mathbb{E}[\Delta W]$. In the lower triangular regions of the heatmap (large absolute delays), overlap is insufficient and learning fails entirely.
5.3 Biological Timescales
Table 1: Biophysical timescales in continuous-time network models. Time constants correspond to distinct molecular/cellular mechanisms: rapid receptor/channel kinetics ($\tau_{\mathrm{prop}}{}$), seconds-scale intracellular signaling and receptor trafficking ($\tau_{\mathrm{plas}}$), and slow homeostatic regulation ($\tau_{\mathrm{dec}}$).
Corresponding Biophysical Process Biological Plausible Range Successful Range (Fig 7)
$\tau_{\mathrm{prop}}{}$ Fast synaptic transmission determined by receptor/channel kinetics and membrane RC filtering such as AMPA/GABAA receptor conductance decay after vesicular glutamate/GABA release (Destexhe et al., 1998; O’Brien et al., 1998). 223030ms in cortex (Kapur et al., 1997); sub-ms in auditory brainstem synapses (Geiger et al., 1997). Swept over 5–30ms, best learning with $\tau_{\mathrm{prop}}{}<20$ms.
[…]
$\tau_{\mathrm{dec}}$ Slow synaptic weakening via protein turnover, phosphatase activity, and homeostatic scaling (Turrigiano et al., 1998; Ehlers, 2003) (e.g., AMPAR endocytosis (Ehlers, 2000), transcriptional regulation). Minutes to tens of minutes. $\sim 20$ minutes. We find that as long as $\tau_{\mathrm{dec}}$ is sufficiently larger than $\tau_{\mathrm{prop}}{},\tau_{\mathrm{plas}}$, it plays little role in the learning dynamics.
A central result of our study is that robust learning requires plasticity windows that outlast the stimulus duration by at least an order of magnitude, placing $\tau_{\mathrm{plas}}$ firmly in the few-second range. This prediction is biologically plausible and experimentally testable, and it narrows the functional range of $\tau_{\mathrm{plas}}$, which has so far been the least constrained timescale in biology.
This requirement emerges when we parameterize our continuous-time networks with biologically motivated constants and evaluate their performance. Real neural circuits operate on nested timescales, from millisecond synaptic conductances to second-scale plasticity windows and slower homeostatic processes. Yet most algorithmic work abstracts away these temporal constraints. By contrast, our simulations demonstrate that the hierarchy $\tau_{\mathrm{prop}}{}\ll\tau_{\mathrm{plas}}\ll\tau_{\mathrm{dec}}$, characteristic of cortical tissue, is sufficient to support effective learning.
In our model, $\tau_{\mathrm{prop}}{}$ corresponds to the dominant synaptic conductance time constant, $\tau_{\mathrm{plas}}$ to the biochemical induction gate during which coincident presynaptic drive and modulatory/error input can trigger plasticity, and $\tau_{\mathrm{dec}}$ to slow synaptic weakening. Once $\tau_{\mathrm{dec}}\gg\tau_{\mathrm{prop}}{},\tau_{\mathrm{plas}}$, it primarily sets a slow baseline for weight decay rather than shaping learning dynamics. Table 1 summarizes these biophysical processes.
Figure 7: Evaluation of a layerwise error routing network on the $7\times 7$ MNIST dataset. Left: 1 hidden layer (49 neurons). Right: 2 hidden layers (49 and 32 neurons). Learning is unstable when the plasticity timescale $\tau_{\mathrm{plas}}$ is comparable to the presentation window ($T=50$ ms), but becomes robust only once $\tau_{\mathrm{plas}}$ exceeds $\sim 2$ s. This corresponds to $\tau_{\mathrm{plas}}/T\approx 40$, highlighting the requirement that plasticity persist far longer than the input presentation time.
Figure 7 evaluates the trained accuracy of two layerwise error routing networks on these biologically grounded timescales. Learning is unstable when $\tau_{\mathrm{plas}}$ is comparable to the input window ($T=50$ ms), but stabilizes once $\tau_{\mathrm{plas}}\gtrsim 2$ s, corresponding to $\tau_{\mathrm{plas}}/T\approx 40$. This finding directly links biological timescales to algorithmic viability, identifying seconds-scale eligibility traces as a necessary condition for feedback-driven learning in cortical-like circuits.
6 Discussion
Our work set out to ask whether biologically plausible learning algorithms can function in continuous time, without relying on artificial phase separation, and at realistic neural timescales. By casting inference and learning as systems of coupled ODEs, we showed that rules such as SGD, FA, DFA, and KP emerge naturally as limiting cases, and that robust learning depends on the temporal overlap of input and error signals. Our simulations further demonstrate that biologically measured timescales are sufficient to support stable learning, leading to a concrete prediction that plasticity windows must persist on the order of seconds.
Limitations. Our empirical evaluation focuses on small datasets and simple tasks due to computational constraints: training of these systems requires explicit forward integration over long time periods of high-dimensional, coupled ODEs whose timescales span over 6 orders of magnitude. This is computationally intensive and numerically delicate (high stiffness, tight tolerances, long horizons), making large benchmarks prohibitive without extensive engineering. Additionally, our neurons are rate-based and simplified, and many biophysical phenomena (spikes, neuromodulatory heterogeneity, etc.) are not modeled. Despite these simplifications, the success of our continuous-time, overlap-driven networks is promising, and future work may address incorporating more faithful biophysics.
Temporal overlap as the operative constraint. In our ODE formulation, the expected update of a synapse is proportional to the temporal cross-correlation between its presynaptic drive $\mathbf{x}$ and a locally projected error $V_{l}^{\top}\bm{\epsilon}_{l}$, low-pass filtered by the plasticity kernel with time constant $\tau_{\mathrm{plas}}$. For piecewise-constant presentations, this yields the analytic overlap kernel $(T-|\Delta|)_{+}$. Failure bands seen at $|\Delta|\approx T$ confirm this prediction: deeper networks accumulate propagation delays that reduce overlap in early layers, degrading learning. This perspective unifies algorithms like FA, DFA, and KP under a single causal principle: a synapse can learn if and only if temporally overlapping input and error reach it. Weight symmetry or fixed feedback are not essential; overlap is.
Plasticity timescale requirement. From the overlap kernel analysis in Section 4, robust learning requires the plasticity kernel to be effectively flat over a presentation window $T$. Allowing at most a fractional attenuation $\eta$ across the window yields $\tau_{\mathrm{plas}}\geq T/\ln\!\big(1/(1-\eta)\big).$ Empirically we observe a transition near $\tau_{\mathrm{plas}}/T\approx 40$, i.e., $\eta\approx 1-e^{-1/40}\approx 0.025$. With $T=50\,\mathrm{ms}$, this places the threshold at $\tau_{\mathrm{plas}}\gtrsim 2\,\mathrm{s}$. Thus, in our protocols, seconds-scale eligibility is sufficient and—absent longer-range tags—apparently necessary; deeper networks and finite propagation tighten this bound by further reducing effective overlap.
Depth and architectural shortcuts. Unlike digital networks, where error gradients are computed simultaneously for all layers, continuous-time networks suffer cumulative delay with depth. This makes learning increasingly fragile as networks grow deeper. Biological circuits appear to mitigate this through architectural shortcuts: skip and feedback pathways are implemented as interareal projections that bypass intermediate layers and, in the neocortex, create shortcuts across the cortical hierarchy (Douglas and Martin, 2004). These connections can provide faster routes for both inference and error signals, preserving temporal overlap in early layers. This interpretation links our results to HSP theory, which argues that arbitrary backward pathways can suffice for credit assignment—our analysis suggests that their role is to preserve overlap, not to enforce symmetry.
Timescales as a design principle. We find that biologically measured timescales are sufficient to support robust learning in our framework. The observed hierarchy $\tau_{\mathrm{prop}}{}\ll\tau_{\mathrm{plas}}\ll\tau_{\mathrm{dec}}$ mirrors the organization of cortical tissue: millisecond synaptic currents, second-scale eligibility traces, and slow homeostatic decay. In simulation, this separation creates a broad regime where learning is stable and accurate, even in the presence of delays and noise. These results suggest that the brain’s hierarchy of clocks may be a functional requirement for maintaining gradient-aligned learning in continuous time.
Implications for hardware. Analog or neuromorphic implementations face the same challenges as biology: finite propagation speeds, integration imperfections, and heterogeneous device dynamics. Our results indicate that hardware does not need to enforce strict weight symmetry, but must preserve overlap between input and error signals across relevant timescales. Designing circuits with fast signal propagation, intermediate eligibility accumulation, and slow decay may therefore be a practical recipe for stable analog learning systems.
Broader perspective. Our continuous-time formulation gives a model for biological and physical learning where inference and plasticity unfold together, constrained by finite propagation and integration dynamics. In this setting, the essential ingredient is not weight transport or symmetry, but the preservation of temporal overlap between inputs and errors within a hierarchy of timescales.
This perspective yields concrete predictions for biology. Deeper circuits are predicted to be more fragile to propagation delays, motivating anatomical shortcuts (skip and feedback pathways) that preserve overlap in early layers. More importantly, robust learning requires plasticity dynamics that outlast the stimulus window by at least an order of magnitude; with cortical integration times of tens of milliseconds, this places the effective eligibility regime in the seconds range. Thus, we predict that eligibility traces supporting error-driven learning must persist at least on the order of seconds—a regime that experimentalists can directly probe by manipulating trace duration or disrupting long-range feedback pathways.
More broadly, our analysis links cortical physiology, algorithmic proposals like FA/DFA/KP, and the practical requirements of analog hardware. It points to a unifying condition—learning succeeds when input and error are temporally correlated at each synapse—that provides both a mechanistic hypothesis for neuroscience and a design rule for neuromorphic systems.
Appendix A Dynamics of a Single Neuron
The single-neuron version of the dynamics presented in equation 1:
$\dot{z}=\frac{-\,z+\sigma(\mathbf{w}^{\top}\mathbf{x})}{\tau_{\mathrm{prop}}},\quad\dot{\mathbf{w}}=-\frac{\mathbf{w}}{\tau_{\mathrm{dec}}^{W}}+\frac{(\mathbf{v}^{\top}\bm{\epsilon})\,\mathbf{x}}{\tau_{\mathrm{plas}}^{W}},\quad\dot{\mathbf{v}}=-\frac{\mathbf{v}}{\tau_{\mathrm{dec}}^{V}}+\frac{(\mathbf{w}^{\top}\mathbf{x})\,\bm{\epsilon}}{\tau_{\mathrm{plas}}^{V}}.$
Appendix B Datasets & Methodology
We use two datasets in our evaluation of these models. The $7\times 7$ downsampled MNIST dataset is the standard MNIST (LeCun, 1998) dataset that has been downsampled with $4\times 4$ average pooling. The circles dataset is the make_circles dataset from scikit-learn (Pedregosa et al., 2011).
[…]
Appendix C Comparison to SGD
To ensure a fair comparison, we train baseline multi-layer perceptrons (MLPs) with the same hidden unit counts as our continuous-time networks with layerwise error propagation. Both models are exposed to the same training data samples in the same order. For the baselines, we use the Adam optimizer (Kingma, 2014) with an initial learning rate of $0.001$, rather than vanilla SGD, to avoid penalizing the baseline with suboptimal hyperparameters. Training is performed with a batch size of 1, matching the presentation schedule used by the continuous-time networks. All networks use ReLU activations (Glorot et al., 2011) on hidden layers.
Table 2 shows accuracies on the $7\times 7$ MNIST dataset. In shallow networks, our layerwise error propagation network matches the baseline on both training and test accuracy, but in deeper networks lags slightly. On simple and fairly linear tasks, this is expected: signal lag accumulates with depth (Section 4), and the dataset itself yields diminishing returns from additional layers.
Train Accuracy Test Accuracy
Model Discrete-time Continuous-time Discrete-time Continuous-time
1-layer $0.8750\pm 0.0054$ $\mathbf{0.9318\pm 0.0028}$ $0.8634\pm 0.0056$ $\mathbf{0.9209\pm 0.0011}$
2-layer $0.8758\pm 0.0220$ $\mathbf{0.9318\pm 0.0032}$ $0.8673\pm 0.0259$ $\mathbf{0.8919\pm 0.0050}$
Table 2: Accuracy (mean $\pm$ std) on the $7\times 7$ MNIST dataset across 3 runs. We compare discrete-time MLP baselines (SGD with Adam, lr=0.001=0.001) to our continuous-time layerwise error propagation network across architectures with 1–2 hidden layers.
On the more nonlinear task of distinguishing concentric circles, additional depth provides a benefit (Table 3). Our continuous-time layerwise error propagation networks achieve comparable or higher test accuracies than the discrete-time baselines.
Train Accuracy Test Accuracy
Model Discrete-time Continuous-time Discrete-time Continuous-time
1-layer $0.8318\pm 0.0276$ $\mathbf{0.9241\pm 0.0133}$ $0.8272\pm 0.0282$ $\mathbf{0.9470\pm 0.0101}$
2-layer $0.9520\pm 0.0076$ $\mathbf{0.9657\pm 0.0082}$ $0.9485\pm 0.0066$ $\mathbf{0.9683\pm 0.0109}$
3-layer $0.9322\pm 0.0300$ $\mathbf{0.9494\pm 0.0588}$ $0.9312\pm 0.0218$ $\mathbf{0.9568\pm 0.0452}$
Table 3: Accuracy (mean $\pm$ std) on the concentric circles dataset across 3 runs. We compare discrete-time MLP baselines (SGD with Adam, lr =0.001=0.001) to our continuous-time layerwise error propagation network across architectures with 1–3 hidden layers.
We can also compare the resilience of our models versus standard SGD in the presence of label error. In our continuous time networks, this arises due to delays in the error signal, as discussed in Section 4. We can model the same error in standard discrete-time neural networks by presenting a standard neural network optimizer a corresponding fraction of mislabeled data. For example, a delay of half the sample time, $\Delta=\frac{1}{2}T$ in our continuous-time model, corresponds to presenting the correct data pairing $(X_{i},Y_{i})$ and then the mismatched label $(X_{i},Y_{i+1})$, where $i$ is the index into the input and label datasets.
Figure 8 compares our continuous-time neural network with a standard discrete-time neural network trained with label dithering, which models the effect of delayed error signals without duplicating training data. Instead of splitting each sample into sub-intervals, we approximate the delay ratio $r=\tau_{\text{delay}}/T_{\text{sample}}\,$ by a rational fraction $p/q$, and then assign exactly $p$ out of every $q$ samples to use the previous label while the rest use the current label. This ensures that the fraction of “mismatched” updates matches the physical delay ratio in expectation, while keeping each sample presented only once. In this way, the training dynamics reflect the temporal overlap of input and delayed error, but the dataset size and number of SGD steps remain unchanged.
Figure 8: Delay ratio versus accuracy. Panels (a), (b), and (c) correspond to networks with 1, 2, and 3 hidden layers. Our continuous-time neural networks with the layerwise error propagation topology demonstrate comparable, if not slightly better, robustness to mismatch than corresponding discrete-time neural networks.